### v2

# Geometric Neural Network based on Phase Space for BCI-EEG decoding

Abstract
Objective: The integration of Deep Learning (DL) algorithms on brain signal analysis is still in its nascent stages compared to their success in fields like Computer Vision, especially in Brain-Computer Interface (BCI), where the brain activity is decoded to control external devices without requiring muscle control. Electroencephalography (EEG) is a widely adopted choice for designing BCI systems due to its non-invasive and cost-effective nature and excellent temporal resolution. Still, it comes at the expense of limited training data, poor signal-to-noise, and a large variability across and within-subject recordings. Finally, setting up a BCI system with many electrodes takes a long time, hindering the widespread adoption of reliable DL architectures in BCIs outside research laboratories. To improve adoption, we need to improve user comfort using, for instance, reliable algorithms that operate with few electrodes. Approach: Our research aims to develop a DL algorithm that delivers effective results with a limited number of electrodes. Taking advantage of the Augmented Covariance Method with SPDNet, we propose the Phase-SPDNet architecture and analyze its performance and computational impact, as well as the interpretability of the results. The evaluation is conducted on 5-fold cross-validation, using only three electrodes positioned above the Motor Cortex. The methodology was tested on nearly 100 subjects from several open-source datasets using the Mother Of All BCI Benchmark (MOABB) framework. Main results: The results of our Phase-SPDNet demonstrate that the augmented approach combined with the SPDNet significantly outperforms all the current state-of-the-art DL architecture in MI decoding. Significance: This new architecture is explainable, with a low number of trainable parameters and a reduced carbon footprint.
Keywords Brain-Computer Interfaces  $\cdot$ Electroencephalography  $\cdot$ Functional connectivity  $\cdot$ SPD manifold  $\cdot$ Riemannian optimization  $\cdot$ Neural Network  $\cdot$ Motor Imagery
1 Introduction
Brain-computer interface (BCI) technology allows direct communication between a user’s brain activity and external devices. Originally designed to help people with disabilities [1], its uses are expanding to other fields, such as rehabilitation using virtual reality [2, 3]. Different signals acquired from brain activity can be used for such a technology, but Electroencephalography (EEG) is a widely adopted choice, as it is a non-invasive, portable, and inexpensive methodology with a very good temporal resolution. Motor imagery (MI) tasks are largely investigated among the brain activities considered when designing a BCI. As the subject is asked to mentally execute a movement without actually performing it, it provides an asynchronous and internal control scheme with no requirement for muscle capability.
The application of Deep Learning (DL) algorithms has garnered significant attention over several domains, ranging from natural language process [4] to protein structure prediction [5]. The field of EEG MI classification is no exception. However, DL algorithms have not yet established themselves in the BCI field as they have in other fields, such as Computer Vision. This is due to several problems: limited data availability, low signal-to-noise ratio in EEG signals, subject variability due to anatomical differences between individuals, and to session variability due to deviations in electrode placement [6].
Unlike other fields where DL algorithms thrive on extensive data, in EEG applications, the emphasis is on enhancing user comfort, resulting in smaller datasets with few trials and fewer electrodes. Large-scale EEG systems with numerous electrodes not only require extended calibration periods, causing user tiredness but also introduce complexity, potentially leading to increased error rates and heightened computational demands. Additionally, the increased cost associated with extensive electrode setups deters the widespread development, deployment, and accessibility of BCI technology, particularly for the patients who could benefit the most from it.
This research focuses on developing a novel DL architecture, Phase-SPDNet , that outperforms the state-of-the-art classification when using a limited number of electrodes. Building on the Augmented Covariance Method (ACM) [7] that is an extension of the spatial covariance and the Symmetric Positive Definite (SPD) Neural Network -SPDNet [8], we study the Phase-SPDNet impact on the performance when using a reduced number of electrodes. We also conducted a comprehensive analysis of the model size, carbon emissions, and, ultimately, its explainability with respect to the standard SPDNet.
Spatial covariance, however, is not the only one that can be extracted from the EEG signal. Another potential candidate is coherence, which provides an alternative perspective on EEG signal characteristics. Historically, coherence has proven to be an inherently unstable feature, which is challenging to compute accurately, resulting in less robust results with respect to the spatial covariance. However, the use of information from both covariance and coherence has been shown to increase performance as we are considering an estimator of the interactions between brain areas/electrodes [9]. We are thus interested to study coherence in combination with SPDNet and especially to use the ACM methodology adapted for the coherence matrix.
Our methodology is tested through a 5-fold cross-validation evaluation (Within-Session), utilizing only three electrodes strategically positioned above the Motor Cortex. To validate our algorithm, we test our approach on almost 100 subjects from openly available datasets using the Mother Of All BCI Benchmark (MOABB) framework [10]. This research not only contributes to the advancement of EEG MI classification but also emphasizes the importance of developing efficient, user-friendly algorithms with a minimal environmental impact in the broader context of BCI technology. Additionally, our study places a strong emphasis on ensuring that our method is both reproducible and interpretable.
The article is organized as follows: In  section 2, we examine the current state-of-the-art in Deep Learning (DL) EEG decoding, emphasizing distinctions from our approach. section 3 provides an overview of the theoretical foundations of our model and the considered datasets. The obtained results from the Within-session evaluation are presented in  section 4. Subsequently, section 5 focuses on the method’s impact and current limitations, specifically focusing on its explainability and environmental impact. Finally, section 6 summarizes the findings of our study.
2 Related Work
2.1 Machine Learning for EEG Decoding
In the domain of EEG decoding [11, 12], translating brain activity into meaningful data has become increasingly dependent on machine learning (ML) methods [13, 14, 6, 15, 9, 16, 17]. However, here, we diverge by focusing on deep learning techniques instead of a broad ML spectrum, moving beyond mere method comparison to introduce a novel EEG decoding approach.
Schirrmeister et al. [14] demonstrated that Deep Learning (DL) approaches, specifically ShallowNet and DeepNet can perform on par with conventional machine learning in decoding raw EEG data and can be trained end-to-end, eliminating several feature extraction steps. The application of DL algorithms enhances the generalization capability, enabling it to handle the inherent variability present in EEG signals effectively. This applicability of DL for EEG data is further corroborated by [18, 19, 20, 21, 15]. Despite their effectiveness, these DL approaches demand extensive parameter tuning, high power consumption, and large datasets for thorough evaluation. In contrast, our study proposes a streamlined approach, leveraging Riemannian geometry and functional connectivity derivatives, which are less demanding in terms of parameters, enhancing both efficiency and scalability. The efficiency and scalability stems from the geometric transformation these methods apply, which reduces the temporal signal’s dimensionality from channel and time dimensions to channel-by-channel dimensions.
2.2 Deep Riemannian Networks for EEG Decoding
Incorporating non-euclidean geometry, especially the Riemannian manifold, into EEG decoding has significantly advanced the field [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. A common practice is to compute spatial covariance matrices that capture signal features with the structure of symmetric positive definite (SPD) matrices. These matrices not only enhance signal information regarding topology and amplitude but also offer increased robustness to outliers and noise with Riemannian geometry, maintaining invariance under linear transformations [24, 29].
Huang and Van Gool [22] work introduced SPDNet, a neural network that operates on SPD manifolds. This foundation was used by Ju et al. [35], who applied SPDNet in bio-signal classification, enhancing transfer learning. Building on this, Brooks et al. [36] adapted batch normalization for the SPD manifold, and Kobler et al. [25] further refined the batch normalization component for EEG Decoding. Pan et al. [24] added to this perspective by proposing attention components for the category of the SPDNet neural networks. Follow-up efforts included using residual layers [26, 31], filter bank inputs [23, 27, 30], mixing traditional convolution by channels [34], and constraining diffusion models [37, 38]. Contrary to our approach, these methods rely on adapting existing neural network components for the SPD manifold or preserving the dimensionality assumptions, while our method combines the dimensionality expansion of the phase space reconstruction on the SPD manifold.
2.3 Geometry Transformation in EEG data
Geometry Transformation (GT) [6, 15, 39, 7] in EEG analysis involves standard pre-processing transformations like resampling, band-passing, and filtering, which have been proven to enhance model performance and biomarker discovery [40, 41, 42, 43, 16]. Data augmentation, a regularization geometry transformation, has been acknowledged for its role in enhancing brain decoding in EEG, with its benefits varying depending on the task [44, 15]. While these transformations can improve the learning processing, these steps usually do not consider the non-stationary dynamics properties of the biological signal. Our approach broadens this scope by incorporating transformations that consider phase reconstruction in terms of its components of nonlinear dynamics.
In the context of delay embedding phase geometry transformation, Chen et al. [45] demonstrated the effectiveness of phase embedding in neural networks for one-dimensional bio-signal reconstruction, but without addressing the non-euclidean nature of the signal. Our methodology, in contrast, considers the non-euclidean geometry, thereby enriching the learning process. The studies by Carrara and Papadopoulo [7] and Zhou et al. [46] align with our approach in their use of phase delays and (cross-)covariance matrices for EEG decoding. Our work, on the other hand, differentiates by fully leveraging the capabilities of the SPD matrix in Riemannian Neural Networks, allowing for more efficient, interpretable layers and achieving superior performance within challenging scenarios with a reduced number of channels.
2.4 Functional Connectivity for EEG Decoding
Beyond the framework of processing the raw brain signal data, researchers in the neuroscience community have been interested in identifying the relationship between brain areas using Functional Connectivity (FC) features [47, 9]. These features demonstrate that electrode interactions are effective for BCI tasks [9, 47, 48, 49, 50, 51, 52, 53, 54]. Most EEG decoding studies using FC employ brain interaction estimators to construct adjacency matrices for each trial, such as covariance, correlation, and coherence, often combined with dimension reduction techniques. We refer to [47] for an extensive discussion on FC estimators for brain signals.
3 Material and Methods
In this section, we describe the EEG decoding problem, the phase space reconstruction transformation, the Riemannian manifold properties, the Symmetric Positive Definite neural network components, and the datasets and baselines.
3.1 EEG Decoding
We consider EEG signals as real-valued matrices. Let us denote each bio-signal window captured during a cognitive task (called an epoch) by $\mathbf{X}_{i}\in\mathbb{R}^{C\times T}$. Here, $C$ represents the number of channels (electrodes), $T$ is the number of time points in the window, and $i$ indexes the trial $i=1\ldots N$, where $N$ is the total number of trials. Each epoch corresponds to a specific cognitive task denoted by $y_{i}\in\mathcal{Y}$.
EEG decoding aims to construct a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, which effectively maps each trial $\mathbf{X}_{i}$ to its corresponding label $y_{i}$. We can construct a neural network function represented as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, composing a sequence of $l$ functions, formulated as $f_{\theta}=f_{\theta_{1}}\circ f_{\theta_{2}}\circ\dots\circ f_{\theta_{l}}$. In this structure, $\theta$ indicates the parameters describing each layer of the network.
3.2 Phase space reconstruction for the EEG signal
When using a reduced number of electrodes, the captured signals will contain only part of the real dynamics of the brain. It is possible to recover part of this information using time delay embeddings based on the Takens’s theorem [55]. This methodology allows the reconstruction of the dynamical system in an alternative space, different from the sensors one, but containing the same dynamical information as the brain.
Our method applies Takens’ theorem over the minimally pre-processed epoch-cropped time series, which allows for the understanding of a system’s multi-variable dynamics through a single observable in a $\psi$-dimensional space by employing a phase embedding transformation as the initial function in our neural network [55, 56, 57]. This is made by constructing a phase space using a delay vector constructed from the original signal, thereby enabling the reconstruction of the system dynamics from these observations. Following the approach proposed by Carrara and Papadopoulo [7], it is possible to define a delay function $d_{p}(\mathbf{\mathbf{X}})=\mathbf{X}(t+p\tau,t+T-(\psi-p)\tau)$ for $p\in[0,\psi]$. This function allows the concatenation of measures on a sliding window over the observable in a delay vector defined as:
$f_{\mathrm{delay}}(\mathbf{X})=[d_{0}(\mathbf{X}),\ldots,d_{\psi}(\mathbf{X})]^{\top}$ (1)
where $\psi$ represents the embedding dimension, which dictates the order of magnitude of the phase space, and $\tau$ is the embedding delay, determining the temporal resolution of our analysis [58, 59]. As a result, $f_{\mathrm{delay}}(\mathbf{X})$ embodies an embedding of the original phase space into a higher-dimensional space, enabling a detailed examination of the system’s dynamics. This formulation considers the EEG signal as produced by a nonlinear dynamical system. Concerning the dimension, if the original epoch have dimension $\mathbf{X}_{i}\in\mathbb{R}^{C\times T}$ the matrix $f_{\mathrm{delay}}(\mathbf{X})\in\mathbb{R}^{C\psi\times(T-\psi\tau)}$. This function is applied on each window signal $\mathbf{X}_{i}\in\mathcal{X}$. An illustration of the process, with $\tau=2$ and $\psi=3$ is present in  Figure 1
Figure 1: Representation of the augmented procedure. The original signal comprises three electrodes, represented by the plot in blue, while the red plot represents the augmented dataset using embedding parameters $\tau=2$ and $\psi=3$.
Selecting the optimal $\tau$ and $\psi$ values can be extremely time-consuming [7, 59]. To overcome this difficulty, we use an adaption of Maximizing Derivatives On Projection (MDOP) [60, 7] method for the multi-epoch context of EEG Decoding, shown in Algorithm 1. The algorithm relies on the function $mdop\_embedding$ from the package $\mathrm{DelayEmbeddings}$ to implement the MDOP procedure. The function aims to create a nonuniform embedding that generates a vector of different lags $\bm{\tau}_{i}$.
Algorithm 1 MDOP for EEG Decoding
using DelayEmbeddings: mdop_embedding
2:function MDOP for epochs($X$)
     $\tau=0$
4:     $\psi=0$
     for $i=1:\#\text{epochs}$ do
6:         $\bm{\tau}_{i}=\text{mdop\_embedding}(X_{i})$
         $\tau=\tau+\lfloor\overline{\bm{\tau}_{i}}\rfloor$
8:         $\psi=\psi+|\text{len}(\bm{\tau}_{i})|$
     end for
10:     $\overline{\tau}_{Epoch}=\lfloor\frac{\tau}{|\#\text{epochs}|}\rfloor$
     $\overline{\psi}_{Epoch}=\lfloor\frac{\psi}{|\#\text{epochs}|}\rfloor$
12:return $\overline{\tau}_{Epoch},\overline{\psi}_{Epoch}$
end function
The MDOP algorithm emerged as a novel geometric method for analyzing dynamical systems, as opposed to conventional statistical and information-theoretic methods, such as mutual information and continuity statistics. Central to MDOP is optimizing an embedding to ensure the reconstructed attractor is as expanded as possible while simultaneously reducing redundancy among the delay components. The algorithm employs a recursive approach, wherein each step of the embedding cycle identifies the lag $\tau$ that yields the highest beta statistics. This chosen lag is then utilized in the subsequent reconstruction phase. This iterative process continues until the algorithm obtains satisfying embedding dimensions $\psi$, selected using the method of false nearest neighbors.
3.3 Riemannian Manifold
Symmetric Positive Definite (SPD) matrices have begun to play a key role in several applications, ranging from brain imaging to Computer Vision [61, 62]. In particular, the classification approach of SPD matrices based on the Riemannian distance algorithm is the current state-of-the-art in the BCI-MI classification [63, 43].
Let’s define $\mathcal{M}_{n}$ the space of real square matrix and $\mathcal{S}^{n}$ the space of symmetric matrix, where $\mathcal{S}^{n}=\{\mathcal{S}^{n}\in\mathcal{M}_{n}\mid\mathbf{S}^{\top}=\mathbf{S}\}$. It is now possible to define the space of Symmetric Positive Definite (SPD) matrices as
$\mathcal{S}_{++}^{n}=\{S\in\mathcal{S}^{n}\mid\bm{x}^{\top}\bm{S}\bm{x}>0\;\;\;\;\forall\bm{x}\in\mathbb{R}^{n}\}$ (2)
This formulation allows us to represent matrices belonging to $\mathcal{S}_{++}^{n}$ as points on a Riemann manifold with a dimension of $n(n+1)/2$. The space of SPD matrices forms a manifold with negative curvature [64, 65], so Euclidean geometry concepts do not apply.
It is possible to define several distances between two SPD matrices $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, generally depending on the length of the geodesic connecting $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ on the Riemann manifold. Often, the affine-invariant metric is used in the context of BCI. We proceed, however, to give a mathematical formulation based on the Log-Euclidean metric [66]. This formulation reduces the computational burden associated with the affine-invariant framework while preserving robust theoretical properties [66].
The space of Symmetric Positive Definite (SPD) matrices can be endowed with a Lie group structure. For a comprehensive understanding, please refer to [66]. It becomes possible to define a distance metric between two SPD matrices $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ as the bi-variate metric on the Lie group of SPD matrices
$d(\mathcal{S}_{1},\mathcal{S}_{2})=||\log(\mathcal{S}_{2})-\log(\mathcal{S}_{1})||$ (3)
where |||||$\quad$||| | | | is the norm associated with the metric, and $\log$ is the matrix logarithm. The Log-Euclidean metric on the Lie group of SPD matrices corresponds to an Euclidean metric within the logarithmic domain of the SPD matrices.
Distances, geodesics, and Riemannian means exhibit a more straightforward formulation within the Log-Euclidean metric than the affine-invariant case, maintaining comparable invariance properties. However, this simplification comes at the cost of more intricate formulations for exponential and logarithmic mapping. The mapping of $\mathcal{S}_{2}$ respect to $\mathcal{S}_{1}$ is defined as
$\begin{array}[]{l}\log_{\mathcal{S}_{1}}\left(\mathcal{S}_{2}\right)=D_{\log\left(\mathcal{S}_{1}\right)}\exp\cdot\left(\log\left(\mathcal{S}_{2}\right)-\log\left(\mathcal{S}_{1}\right)\right),\\ \exp_{\mathcal{S}_{1}}(\mathcal{S}_{2})=\exp\left(\log\left(\mathcal{S}_{1}\right)+D_{\mathcal{S}_{1}}\log\cdot\mathcal{S}_{2}\right).\end{array}$ (4)
where $D_{\log\left(\mathcal{S}_{1}\right)}\exp$ represent the differential at point $\mathcal{S}_{1}$ of the exponential function and similarly for $(D_{\log\left(\mathcal{S}_{1}\right)}\exp)^{-1}=D_{\mathcal{S}_{1}}\log$. This formulation can be simplified if the reference matrix for the mapping is the identity matrix. Equation 4 presents two quantities, the exponential and logarithmic maps, which are used to transition from Euclidean space to the Riemannian surface.
In BCI, the spatial covariance is estimated from the pre-processed EEG signal $\mathbf{X}\in R^{C\times T}$. Several estimators can be used to estimate covariance [67], but the most popular is the sample covariance matrix
$f_{\mathbb{C}\text{ov}}(\mathbf{X}_{i})=\frac{1}{T-1}\sum_{i=1}^{T}\mathbf{X}_{i}\mathbf{X}_{i}^{\top}$ (5)
3.3.1 Symmetric Positive Definite Neural Networks Components
In the context of geometry neural networks, Huang and Van Gool [8] proposed the neural network SPDNet with three SPD layers:
BiMap Layer:
The bi-linear mapping level aims at creating more compact and discriminating SPD matrices just like the convolutional level, with the complication, however, that SPD matrices live in a Riemannian space. We can describe the equation as
$f_{\mathrm{BiMap}}(\mathbf{Z}_{k-1})=\mathbf{Z}_{k}=\textbf{W}_{k}\mathbf{Z}_{k-1}\textbf{W}_{k}^{\top}\;,$ (6)
where $\mathbf{Z}_{k-1}\in\mathcal{S}_{d_{k-1}}^{++}$ and W is, for this layer, the learnable parameter. Because the output $\mathbf{Z}_{k}$ must belong to the SPD space, we require that $\textbf{W}_{k}\in\mathbb{R}^{(d_{k},d_{k-1})}$ belongs to a compact Stiefel $St(d_{k},d_{k-1})$ manifold.
ReEig Layer:
This layer introduces a non-linearity with a similar approach to a Rectified Linear Unit (ReLU) level. In practice, it rectifies SPD matrices by thresholding small eigenvalues to $\varepsilon$.
$f_{\mathrm{ReEig}}(\mathbf{Z}_{k-1})=\mathbf{Z}_{k}=\textbf{U}_{k-1}\max(\varepsilon I,\Sigma_{k-1})\textbf{U}_{k-1}^{\top}\;,$ (7)
where $\textbf{U}_{k-1}$ and $\Sigma_{k-1}$ are not learned but obtained using eigenvalue decomposition of the previous layer, $\mathbf{Z}_{k-1}=U_{k-1}\Sigma_{k-1}U_{k-1}^{T}$. This layer does not present any trainable parameter.
LogEig:
This layer aims to transport the SPD matrix obtained from the previous layers from a Riemannian space to an Euclidean one using the Log-Euclidean metric [66]. This is formally the expression (4) considering the mapping with respect to the identity, so formally, the layer is defined as
$f_{\mathrm{LogEig}}(\mathbf{Z}_{k-1})=\mathbf{Z}_{k}=\log(\mathbf{Z}_{k-1})=\textbf{U}_{k-1}\log(\Sigma_{k-1})\textbf{U}_{k-1}^{\top}$ (8)
where again, the $\textbf{U}_{k-1}$ and $\Sigma_{k-1}$ are not learned but obtained using eigenvalue decomposition. This layer not present any trainable parameter. This logarithmic mapping is applied using the Identity as a reference matrix. Once this layer is applied, we can use the classical deep learning method in the Euclidean space as a Multi-Layer Perceptron.
Given the framework, we define our neural network, denoted as $f$, as a composition of sequential transformations:
$f_{\mathrm{Phase-SPDNet}}=f_{\mathrm{delay}}\circ f_{\mathbb{C}\text{ov}}\circ f_{\mathrm{BiMap}}\circ f_{\mathrm{ReEig}}\circ f_{\mathrm{LogEig}}\circ f_{\mathrm{MLP}},$ (9)
where each $f_{l}$ represents a specific transformation or layer within the network. The learning process is formalized as the mapping $f_{\mathrm{Phase-SPDNet}}:\mathcal{X}\rightarrow\mathcal{Y}$, which operates on the training dataset. $\theta$ denotes the set of parameters within the parameter space $\Theta$. The optimization objective is to minimize an average loss $\ell$ over the training dataset, defined as:
$\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}\ell(f_{\mathrm{Phase-SPDNet}}(\mathbf{X}_{i}),\mathbf{y}_{i}).$ (10)
The model’s generalization capability is further assessed using an independent test set.
In a DL typical training process, optimization methods like Adam [68] are normally used for implementing the back-propagation procedure. However, in the context of SPDNet, we must learn the weights of $W$ of the BiMap layer in such a way that the new weights still belong to the Stiefel manifolds, i.e., they are still orthonormal matrices. To solve the problem, we use the RiemannAdam [69] optimization from the geoopt [70] library. In addition to the optimization strategy, we use the standard cross-entropy loss function $\ell$ as a loss function, a well-established and widely used objective function in classification tasks.
To summarize, the resulting architecture Phase-SPDNet uses enriched SPD matrices that encapsulate a broader spectrum of information than their traditional counterparts thanks to the additional layer that implements the phase space reconstruction. In addition, such an approach allows, through the use of nonlinear systems theory, the reconstruction of a phase space that contains more information with respect to the one extracted from the original signal, allowing the use of fewer electrodes. The augmented SPD matrices require an SPDNet with a larger number of parameters, which is partly counteracted by using fewer electrodes. In particular, this representation accentuates how the input SPD matrix is modified from the initial dimension $\mathcal{S}_{C}^{++}$ to $\mathcal{S}_{C\times\psi}^{++}$.
Overall, this enrichment of the information contained in the SPD matrices not only expands the range of discriminative information that can be extracted but also forces the network to adapt to more intricate and fuzzy patterns in the data. Figure 2 provides a graphical picture of our methodology.
Figure 2: Overview of our approach. A. Phase space reconstruction for the EEG signal. For each trial from $\mathbf{X}$, we apply the function $f_{delay}$ to reconstruct the phase space. During each evaluation fold, we adopted the MDOP algorithm to estimate suitable function parameters for embedding dimension $\psi$ and the embedding delay $\tau$. For illustrative purposes, this figure considers 3 electrodes located over the motor cortex with parameters $\psi=2$ and $\tau=10$. B. Conversion of the phase space time-series with $f_{\mathbb{C}\text{ov}}$ for the covariance representation of the Symmetric Positive Definite Space $\mathcal{S}_{++}^{n}$, represented in green. With the parameter considered the covariance is a square matrix of dimension 6. C. The covariance feature space is then used to train the Symmetric Positive Definite Neural Network with BiMap, ReEig, LogEig, and MLP layers. Once the representation returns to the Euclidean space (red), after the LogEig, we adjust a fully connected layer to predict the associate label $y$ for each trial.
In the case of Phase-SPDNet , we set the subspace dimension of the BiMap layer at half of the input dimension [8]. However, for standard SPDNet, we maintained the subspace dimension at the same value as the original input dimension (although in this case the BiMap layer acts only as a rotation) in order to proceed with subsequent analyses concerning the explainability of the model. Anyway, diminishing the subspace dimension of half in the case of SPDNet consistently led to a reduction in performance. Note that the current implementation of SPDNet has a ReEig layer that is not scale-independent because of the parameter $\varepsilon=10^{-4}$. To enhance the architecture’s scale independence, we have introduced a standardization procedure for the raw signal, bringing every channel to a zero mean and a unit standard deviation.
In short, Phase SPDNet processes the SPD matrix that contains the spatial information and the temporal evolution of the signal, while SPDNet handles only the spatial covariance.
3.4 Datasets
In order to assess the replicability of our approach, we used six different open datasets from MOABB [10] consisting of almost 100 subjects as shown in Table 1.
It turns out that it is essential to employ different datasets to obtain reliable results, especially when evaluating different ML techniques in EEG decoding, in order to avoid the bias of a single data set [13].
Dataset Subjects Channels Sampling Rate (Hz) Sessions Tasks Trials/Class Epoch (s)
BNCI2014001 [71] 9 22 250 2 4 144 [2, 6]
BNCI2014004 [72] 9 3 250 5 2 360 [3, 7.5]
Cho2017 [73] 52 64 512 1 2 100 [0, 3]
Schirrmeister2017 [74] 14 128 500 1 4 120 [0, 4]
Weibo2014 [75] 10 60 200 1 7 80 [3, 7]
Zhou2016 [76] 4 14 250 3 3 160 [0, 5]
Table 1: Motor Imagery datasets considered during this study
We employed standard pre-processing steps designed for all datasets described in [10]. We applied band-pass filtering with the overlap-add method between $[8-32]$ Hz [77, 78], artifact rejection, and electrode standardization, bringing every channel to a zero mean and a unit standard deviation. Additionally, we opted to use the complete epoch duration, although this duration varies among datasets, as shown in Table 1. The pre-processing steps just mentioned are the same for all the pipelines considered in our study.
We concentrate on a binary classification task, aiming to discern between imagined movements of the right and left hand using an intra-subject/within-session evaluation, which is based on a $5-$fold cross-validation is conducted in each session.
Additionally, we focus on investigating our algorithm’s robustness with fewer electrode settings. For this purpose, our channel selection process was guided by two constraints: ensuring neurophysiological relevance and maintaining consistent electrode selection across datasets. We selected electrodes positioned over the sensorimotor region, particularly in the central (C) and centro-parietal (CP) lines, which are actively involved in the tasks. This choice is strongly supported by Brain-Computer Interface (BCI) research [78, 79, 80, 81, 82, 83, 84]. For consistency, we used EEG montages with varying numbers of electrodes, from 3 to 128. We standardized our analysis by using the minimal overlapping electrodes common to all montages: C3, Cz, and C4.
3.5 Baseline Comparison
We compare the performance of our pipeline to that of several state-of-the-art DL Neural Networks used for BCI. All algorithms in our study were tested using the same set of three electrodes to maintain uniformity in our experimental approach. To standardize the models’ parameters, we initially resample the input time series to align the signal with the state of the art in order to use the parameter’s architecture of the original paper.
  1. 1.
    ShallowNet [14], a neural network architecture with independent spatial and temporal convolution steps with Relu activation, with standardized and re-sampled EEG signal at 250Hz.
  2. 2.
    DeepNet [14], an approach similar to ShallowNet with a deeper linear layer, with standardized and re-sampled EEG signal at 250Hz.
  3. 3.
    EEGNet [18], have a depth-wise convolutional layer functions as a spatial filter across channels, complemented by a separable convolution layer for features extraction designed for EEG with a sample frequency of 128Hz.
  4. 4.
    EEGTCNet [19], is a neural network that wrapper the EEGNet and includes a Temporal Convolution Network over the embedded representation, with standardized and re-sampled EEG signal at 250Hz.
  5. 5.
    EEGITNet [20], a neural network inspired by InceptionNet with parallel convolution layers with different scales. The network has been designed for EEG signals sampled at 128Hz.
  6. 6.
    EEGNeX [21], a neural network inspired by EEGNet incorporation the key components from ConvNeXt. The network has been designed for EEG signals sampled at 128Hz.
  7. 7.
    DynSpat-EEGNet [85], a neural network based on Dynamic spatial filtering with the EEGNet as head. The network has been designed for EEG signals sampled at 250Hz.
  8. 8.
    DynSpat-ShallowNet [85], a neural network based on Dynamic spatial filtering with the EEGNet as head. The network has been designed for EEG signals sampled at 250Hz.
These methods were implemented and trained using MOABB version 1.0  [10]. For in-depth information about the deep learning hyper-parameters, please consult Table A2. In order to obtain a fair time comparison, all the following results are computed on the same hardware, a Dell C6420 dual-Xeon Cascade Lake SP Gold 6240 @ 2.60GHz.
4 Results
In the following section, we analyze the results produced using within-session (WS) evaluation. This method is based on a 5-fold cross-validation performed on each session independently. In particular, we report two separate analyses.
The first analysis is dedicated to contrasting the performance of Phase-SPDNet , with a specific focus on the exploitation of the covariance feature, against the state-of-the-art techniques in DL applied in EEG. The subsequent analysis, reported in the Appendix, shifts the focus towards the exploration of coherence as a feature for classification. Despite the ongoing investigations into coherence, we deliberately separated this analysis due to its comparatively lesser performance against the benchmarks achieved with covariance features. Nevertheless, the augmentation procedure is a significative methodology to improve the performance of both feature sets.
For the first analysis, we tested the performance of Phase-SPDNet against the baseline comparison defined in 3.5 across a broad spectrum of datasets. The results are listed in Table 2, while the detailed statistical analysis can be found in Figure 3.
Table 2 shows ROC AUC results for right-hand versus left-hand classification, where bold numbers represent the best score in each dataset, revealing a consistent trend wherein our algorithm Phase-SPDNet surpasses the DL state-of-the-art in the context of reduced datasets. This behavior is true for all datasets, with the notable exception of the Schirrmeister2017 dataset, where it is marginally surpassed by ShallowNet, DeepNet and DynSpat-ShallowNet. It is important to note that the results of the first two, however, are not statistically significant for this dataset, as presented in Figure 3 (c), (d). The results obtained in this dataset by DynSpat-ShallowNet is statistically superior, as shown in Figure 3 (e). However, we are interested in computing a statistic behavior considering all datasets under consideration. As the number of subjects per dataset is actually too small to reach a decent statistical power for investigating the differences between pipelines, the results across datasets are aggregated to conduct a meta-analysis. Our Phase-SPDNet yields statistically better results than all other approaches as shown in Fig. 3 (b), where only significant interactions are displayed ($p$ > 0.0.5), the reported values and color show the $t$-values.
The second relevant point is to see the impact of the augmentation procedure compared to the standard SPDNet. Figure 3 (a) elucidates this aspect by illustrating the relative performance enhancements of various models against the SPDNet standard. Notably, ShallowNet and DeepNet exhibit positive performance increments across the majority of the datasets examined. However, the augmentation procedure, as implemented by Phase-SPDNet , not only consistently yields positive relative improvement against the standard SPDNet across all datasets but also the improvement introduced by the augmentation procedure turns out to be so significant that it even outperforms methods that initially surpassed the SPDNet benchmark. This improvement is not only demonstrated by the relative improvement with respect to the SPDNet standard but also is evident through the meta-analysis comparison (refer to Fig 3 (c), (d) and (e)).
Ultimately, the statistical analysis conducted using Figure 3 (b) allows us to demonstrate that Phase-SPDNet outperforms the state-of-the-art DL methodology in a reduced dataset context. Its versatility and reliability, particularly in data-limited scenarios, position the method as a promising solution for real-world applications.
Table 2: Average AUC-ROC (%) on six datasets (BCNI2014001, BCNI2014004, Cho2017, Schirrmeister2017, Weibo2014, and Zhou2016), for a left vs. right motor imagery task. The highest performances are in bold-face.
Models BNCI2014001 BNCI2014004 Cho2017 Schirrmeister2017 Weibo2014 Zhou2016
DeepNet [14] (2017) 75.80 $\pm$ 15.45 72.80 $\pm$ 19.48 63.13 $\pm$ 14.45 73.04 $\pm$ 15.67 73.97 $\pm$ 18.07 91.74 $\pm$ 7.00
ShallowNet [14] (2017) 75.85 $\pm$ 15.41 72.17 $\pm$ 18.61 64.14 $\pm$ 13.03 73.59 $\pm$ 15.19 75.36 $\pm$ 15.69 88.03 $\pm$ 8.55
EEGNet [18] (2018) 70.64 $\pm$ 19.87 70.27 $\pm$ 18.91 60.23 $\pm$ 14.98 69.80 $\pm$ 16.51 71.94 $\pm$ 17.80 88.95 $\pm$ 7.84
EEGTCNet [19] (2020) 65.98 $\pm$ 17.25 66.86 $\pm$ 18.41 56.43 $\pm$ 12.31 67.87 $\pm$ 17.36 65.94 $\pm$ 15.69 81.47 $\pm$ 11.66
EEGITNet [20] (2022) 66.64 $\pm$ 13.98 64.93 $\pm$ 14.49 54.68 $\pm$ 11.97 62.98 $\pm$ 16.19 56.97 $\pm$ 17.66 72.82 $\pm$ 12.78
EEGNeX [21] (2022) 68.86 $\pm$ 17.27 68.29 $\pm$ 17.85 56.64 $\pm$ 12.83 64.26 $\pm$ 17.58 58.73 $\pm$ 19.40 80.25 $\pm$ 15.55
DynSpat+EEGNet [85] (2022) 68.06 $\pm$ 18.58 76.61 $\pm$ 16.81 61.51 $\pm$ 13.74 70.62 $\pm$ 16.53 72.88 $\pm$ 19.33 87.85 $\pm$ 7.01
DynSpat+ShallowNet [85] (2022) 70.95 $\pm$ 18.78 78.95 $\pm$ 17.23 65.52 $\pm$ 14.65 75.07 $\pm$ 14.40 74.70 $\pm$ 19.12 87.33 $\pm$ 8.87
SPDNet [8] (2017) 71.02 $\pm$ 15.64 70.15 $\pm$ 16.85 59.95 $\pm$ 12.61 67.40 $\pm$ 13.01 67.04 $\pm$ 17.62 88.85 $\pm$ 8.05
Phase-SPDNet (Our) 75.98 $\pm$ 17.00 80.46 $\pm$ 16.64 66.00 $\pm$ 13.29 72.02 $\pm$ 13.07 78.01 $\pm$ 19.64 94.92 $\pm$ 3.34
Figure 3: Results for Right vs Left-hand classification, using Within-Session evaluation. Plot (a) provides the relative improvement of the AUC-ROC in percentage of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet against SPDNet, ShallowNet, and DynSpat+ShallowNet, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.
5 Discussion
5.1 Impact of augmentation on most and least responsive 5 subject of Cho2017
In the previous section, we explored the outcome obtained by Phase-SPDNet within the context of a reduced number of electrodes. Now, our attention shifts to a deeper exploration of this new algorithm, with a specific emphasis on the interpretability and computational performance of the method. To facilitate this comprehensive analysis, still in the context of three electrodes, we have chosen to concentrate on the examination of the five most and least responsive subjects within the Cho2017 dataset (the wider dataset in terms of the number of subjects), selected using the Minimum Distance to the Mean (MDM) algorithm [63] with covariance as a feature.
To analyze the performance, we use different pipelines, MDM [63], ACM+MDM [7], SPDNet, and Phase-SPDNet , each offering a distinct perspective on the impact of augmentation techniques on both traditional ML and advanced DL algorithms. Furthermore, we assess the performance using three different SPD features: covariance, instantaneous coherence, and imaginary coherence (Figure 4).
Figure 4: (a) Plot showing the box plot of the five most responsive subjects (3, 14, 35, 41, 43) of Cho2017. We see that the augmentation procedure consistently leads to notable improvements in the classification performance for all the SPD estimators considered.(b) Plot showing the box plot of the least responsive five subjects (2, 7, 29, 34, 50) of Cho2017. We see that the augmentation procedure increases the performance for the covariance feature, while the improvement is less clear for the FC estimator. In the plot, the names Ins and stand respectively for Instantaneous Coherence and Imaginary Coherence.
For the five most responsive subjects, covariance stands out as the top-performing feature as it could be expected. However, in the case of the five least responsive subjects, this dominance of covariance is less clear when considering the state-of-the-art methodology.
The outcomes of our study are in concordance with the state-of-the-art [9]: results achieved using Covariance as SPD feature outperform those obtained with the Functional Connectivity (FC) estimator.
It becomes evident that the augmentation procedure consistently leads to notable improvements in the classification performance for the top-performing five subjects. However, the situation becomes less definitive for the least responsive five subjects, where the augmentation effect is not as pronounced or evident. In this case, their signals exhibit a lower SNR, which can substantially impact the MDOP algorithm, leading to a poor estimation of the hyper-parameter. In this case, the augmentation procedure achieves a very positive impact on covariance while there is even a decrease in performance using coherence.
5.2 Interpretability
Creating explainable algorithms is crucial in the field of DL, and it holds even greater significance in healthcare data. In healthcare, understanding why an algorithm makes specific decisions is of great importance for clinicians and patients, ensuring transparency, accountability, bias mitigation, education, and trust-building in the collaboration between AI systems and medical experts.
For analyzing the explainability of the Phase-SPDNet we use the GradCam++ algorithm [86], a technique designed for visualizing and explaining the decision-making process of neural networks. The GradCam++ output is essentially a visual representation that reveals which elements within the data are the most important for the model. GradCam++ has proven to be a robust method for spatial interpretability of EEG signals in a class-agnostic manner [87].
In our analysis, we focus on Subject 41 from the Cho2017 dataset, which is notable for being among the top 5 subjects that exhibit the most substantial performance improvement through the augmentation procedure. Similar findings also hold for other subjects. In particular, using covariance as a feature, we show a sample of the test dataset that was correctly classified by the Phase-SPDNet algorithm (left hand classified as left hand) but was misclassified by SPDNet using the same feature, as shown in Figure 5. Similar findings have been observed for the FC estimators. This figure represents the output of the GradCam++ applied to the output of the ReEig layer.
In all three scenarios, the standard SPDNet places more emphasis on the diagonal terms of SPD matrices. On the other side, when employing Phase-SPDNet , a more comprehensive picture unfolds. While diagonal terms continue to play a pivotal role, what becomes increasingly evident is the significant contribution of the off-diagonal terms. In essence, this observation emphasizes that the Phase-SPDNet model leverages both individual electrode characteristics (diagonal terms) and the interplay between electrodes in time and space (off-diagonal terms) as crucial elements in its classification decision-making process.
Figure 5: Interpretability parallel between GradCam++ and t-test between right hand left-hand classification for subject 41 of Cho2017. Plot (a) is the GradCam++ visualization, applied after the ReEig layer, for Augmented covariance. Plot (b) is a zoom in the region corresponding to standard Covariance. Plot (c) is the GradCam++ visualization, applied after the ReEig layer for standard covariance. Plot (d) is the t-test of the Augmented covariance matrix. Plot (e) is the t-test of the covariance matrix. The t-test visualization represents only the significant interactions ($p<0.05$). For the standard covariance matrix, the only significant interaction occurs between electrodes C3, which mirrors the representation seen in the GradCam++ visualizations. In contrast, the Phase-SPDNet (Figure 5(d)) reveals additional significant interactions, although this interaction map does not precisely match the GradCam++ visualization. This discrepancy highlights the added value of our method in identifying crucial inter-channel relationships that might be overlooked by traditional covariance analysis.
To investigate deeper into the interpretability of our proposed network, we employ a paired t-test to highlight statistically significant differences in the augmented covariance matrix (ACM) between left and right hands. Our analysis centers on subject 41 from Cho2017, specifically on the test dataset of the initial fold of the cross-validation, with consistent findings observed in the other folds. Given that 20 samples are available for each imagined movement, we are in the domain of application of the t-test.
Examining Figure 5 (d) and (e), it is evident that SPDNet places emphasis primarily on the statistically significant pair of electrodes, showing a clear parallel between the GradCam++ visualization and the t-test. In contrast, Phase-SPDNet not only considers significant interactions but also extends its focus to nearly significant interactions.
5.3 Convergence Behavior
To analyze convergence behavior, we again consider subject 41 of the Cho2017 dataset. Again, to get a fair comparison, all the following results are calculated on the same hardware, a Dell C6420 dual-Xeon Cascade Lake SP Gold 6240 @ 2.60GHz.
We examine the outcomes concerning both training and validation metrics, including loss and Receiver Operating Characteristic Area Under the Curve (ROC AUC) scores. The consolidated results are visually presented in Figure 6, wherein the graphs depict the average values across the considered 5-fold cross-validation.
Notably, our observations reveal that the loss function of Phase-SPDNet converges to zero at a notably accelerated pace and exhibits a considerably more stable trajectory compared to standard SPDNet. Moreover, while the loss function of Phase-SPDNet steadily approaches zero, the standard SPDNet experiences a more gradual descent, deviating from the path to complete convergence but ultimately plateauing around $0.4$. These findings highlight that the capacity of SPDNet is limited and that the algorithm stops learning at a certain point.
Interestingly, when examining the ROC AUC graph, we observe that the validation scores of Phase-SPDNet converge consistently to higher values. Notably, this convergence is achieved even faster compared to the SPDNet standard, requiring only $94\%$ of the iterations. This implies that Phase-SPDNet attains a comparable performance to SPDNet but with a significantly reduced number of iterations, making it an efficient and effective choice for real-life applications.
Figure 6: Convergence behavior for right-hand vs. left-hand classification using WS evaluation. The plot uses subject 41 of the Cho2017 dataset, showing the mean value for each plot computed over the 5-Fold considered in the WS evaluation. Plot (a) provides the training loss, while plot (b) is the validation one. Plot (c) provides the ROC-AUC for the training phase, while plot (d) the one for the validation.
5.4 Computational Performances
In this section, we are interested in the analysis of computational performance, with a specific focus on two key aspects: execution time and environmental impact. The results on the best 5 subjects of Cho2017 (subjects 3, 14, 35, 41, 43) are presented in Table 5.4.
Recently, much attention has begun to be given to the sustainability of ML and AI algorithms [88]. Since more and more models continue to proliferate and demand substantial computational resources, it is crucial to estimate and mitigate the energy consumption and emissions associated with their training and deployment. To assess the environmental impact of our algorithm, we use the CodeCarbon [89] library. CodeCarbon is an open-source library that enables developers to monitor carbon dioxide ($CO_{2}$) emissions, offering a way to analyze the carbon footprint of our work.
Interestingly, our analysis shows that the Phase-SPDNet exhibits the longer computational time among the evaluated algorithms while simultaneously standing out as one of the most environmentally friendly in terms of its $CO_{2}$ impact (see Fig 5.4). This intriguing juxtaposition can be attributed to the fact that, during each fold of the cross-validation process, there’s a necessity to optimize the model’s hyper-parameters, which naturally augments the computational time. However, the MDOP algorithm consumes fewer resources than gradient descent procedures, thus resulting in a comparatively smaller environmental impact. In fact, the use of augmented procedure introduces the complication of two hyper-parameters but nevertheless manages to keep the number of trainable parameters of the model contained (see Table 5.4). It’s plausible that our algorithm has yet to reach its optimal performance. Further investigation is needed to improve the optimization strategies.
[…]
figurePerformance (ROC AUC) versus $CO_{2}$ consumption (g $CO_{2}$ equivalent). The size of the blob indicates the number of trainable parameters.
Avg. Time (s) $CO_{2}$ Emission (g $CO_{2}$)
DeepNet 21.77 0.39
ShallowNet 15.10 0.25
EEGNet-8,2 14.96 0.25
EEGITNet 28.44 0.48
EEGTCNet 28.83 0.47
EEGNeX 28.15 0.46
DynSpat+ShallowNet 34.42 0.41
DynSpat+EEGNet 34.09 0.45
SPDNet 6.80 0.07
Phase-SPDNet 30.16 0.22
[…]
5.5 Limitation and Future Directions
The existing Phase-SPDNet , as employed thus far, utilizes MDOP for hyper-parameter selection, offering computational efficiency compared to grid-search methods [7]. However, the solutions obtained with MDOP were sub-optimal compared with using a more extensive hyper-parameter selection. A potential avenue for future exploration involves streamlining and accelerating the grid-search process to harness the capabilities of the Phase-SPDNet architecture fully.
Further enhancements could be achieved by adopting more intricate SPDNet architectures, such as incorporating batch normalization, exploring different BiMap layers, or experimenting with a Bottleneck architecture.
An intriguing direction arises from the insights gained in the explainability study, revealing that Phase-SPDNet predominantly focuses on diagonal elements. This suggests the possibility of employing a Region of Interest approach, where the network initially prioritizes the main diagonals while treating the remainder of the matrix as background.
Continued improvements could be realized through comprehensive algorithm testing across diverse datasets, tasks, and evaluation procedures. This approach allows for a better understanding of performance in more complex scenarios, particularly those posed by intra- and inter-subject variabilities.
Using three electrodes only is advantageous for practical scenarios. In fact, this decision was strategically made to shorten the length of the experiment, thereby minimizing patient fatigue and maintaining their attention throughout the session.
Another challenge was the difficulty in classifying data from less responsive subjects. Despite our efforts, there remains a need for improved methods better to capture the subjects’ intentions and their specific characteristics.
6 Conclusion
This study focuses on integrating the Augmented Covariance Matrix (ACM) with SPDNet to develop a practical algorithm intended for real-world applications, with a deliberate focus on scenarios involving a limited number of electrodes. Specifically, we restricted our electrode usage to just 3, strategically placed on the motor cortex.
In fact, the augmented methodology, based on Takens theorem, demonstrates a particularly effective performance when applied with a reduced number of electrodes. However, it is important to note that using only three electrodes can impact the results. For instance, in the BNCI2014001 dataset, employing just three electrodes led to a performance reduction of about $10\%$ in the ShallowNet architecture compared to using all available electrodes.
Validation of our approach involves leveraging nearly 100 subjects from several open datasets, a task facilitated by the MOABB framework. The resulting Phase-SPDNet algorithm outperforms existing DL algorithms in BCI-EEG classification and offers some explainability elements through GradCam++ visualization. Remarkably, this algorithm requires a modest number of trainable parameters and exhibits a lower environmental impact, measured by $CO_{2}$ equivalent emissions.
7 Acknowledgments
The work of IG and TG was partly funded by the EUR DS4H/Neuromod fellowship. The BA work was supported partly by CAPES under Grant 001 and by DATAIA Convergence Institute as part of the “Programme d’Investissement d’Avenir”, (ANR-17-CONV-0003) operated by LISN. We extend our gratitude towards the OPAL infrastructure at Université Côte d’Azur for their essential resources and support. Additionally, acknowledgment is due for the support from the European Research Council (ERC) under the EU’s Horizon 2020 research and innovation program (grant No. 864729), and the “Investissements d’avenir” program ANR-10-IAIHU-06.
8 Open-source availability
The code will be made public during publication time under the BSD-3 License.
Appendix A Results using Coherence as feature
In this section, we shift our focus to reporting results due to the use of Imaginary and Instantaneous coherence features as input. Specifically, we aim to draw a comparison between the Phase-SPDNet results and the non DL state-of-the-art pipelines (usually created for covariance but here applied on coherence). It’s worth noting that, in the subsequent analysis, we abstain from comparing these results with the performance of state-of-the-art DL models, given that the employment of coherence-based features typically yields lower results across the board. Nevertheless, we find it particularly intriguing to present the outcomes achieved by Phase-SPDNet in the context of coherence-based features since we observe that Phase-SPDNet demonstrates a statistically significant improvement. This shows the potential and significance of Phase-SPDNet in the broader landscape of signal processing and pattern recognition.
  1. 1.
    CohCSP+LDA [90], a combination of the Common spatial pattern (CSP) algorithm followed by a classification performed on a shrinkage Linear Discriminant Analysis (LDA).
  2. 2.
    Coh + TS + EN [9], a Riemannian classification method in the tangent space using an Elastic Network as classifier.
  3. 3.
    CohFgMDM [91], a Riemannian classification method by Minimum Distance to the Mean after having applied a geodesic filtering.
  4. 4.
    CohMDM [63], a Riemannian classification method by Minimum Distance to the Mean.
  5. 5.
    Coh + TS + SVM [63], a Riemannian classification method in the tangent space using a Support Vector Machine (SVM).
Results can be found in Table A1 while the detailed statistical analysis can be found in Figures A1 and A2.
Table A1: Summary of performances via average AUC-ROC (%) on six datasets (BCNI2014-01, BCNI2014-04, Cho2017, Schirrmeister2017, Weibo2014, and Zhou2016), for a left vs. right motor imagery task. Bold numbers represent the best score in each dataset. Estimators Imaginary Coherence and Instantaneous Coherence
Models Estimator BNCI2014001 BNCI2014004 Cho2017 Schirrmeister2017 Weibo2014 Zhou2016 SPDNet Imn 56.55$\pm$7.41 58.12$\pm$12.20 56.16$\pm$8.94 59.15$\pm$7.84 60.59$\pm$10.78 62.89$\pm$8.97 CSP+LDA Imn 57.66$\pm$7.19 53.09$\pm$8.55 51.63$\pm$6.30 53.06$\pm$5.20 56.77$\pm$7.62 64.47$\pm$7.47 TANG+SVM Imn 59.39$\pm$8.86 53.66$\pm$8.26 52.45$\pm$6.24 52.93$\pm$4.83 56.09$\pm$8.89 64.19$\pm$9.49 FgMDM Imn 60.19$\pm$7.72 52.83$\pm$8.90 51.93$\pm$6.39 53.99$\pm$4.57 57.50$\pm$7.92 65.52$\pm$8.72 MDM Imn 60.30$\pm$7.83 54.04$\pm$9.00 52.09$\pm$6.67 53.36$\pm$5.34 57.31$\pm$7.33 66.27$\pm$8.60 Cov+EN Imn 60.71$\pm$7.92 52.67$\pm$8.96 52.07$\pm$6.70 54.42$\pm$4.07 57.21$\pm$8.23 64.58$\pm$8.41 Phase-SPDNet Imn 60.89$\pm$10.17 65.70$\pm$14.06 56.51$\pm$8.60 61.69$\pm$9.81 64.75$\pm$11.33 72.16$\pm$7.40 SPDNet Ins 63.42$\pm$5.76 63.71$\pm$12.88 57.61$\pm$11.09 59.73$\pm$6.89 66.11$\pm$12.02 75.07$\pm$7.63 MDM Ins 69.74$\pm$14.19 59.89$\pm$10.72 58.76$\pm$9.65 52.01$\pm$5.77 63.24$\pm$14.25 81.94$\pm$8.91 TANG+SVM Ins 69.82$\pm$14.76 60.44$\pm$9.76 58.86$\pm$10.11 53.31$\pm$5.42 63.10$\pm$15.51 85.02$\pm$8.29 CSP+LDA Ins 70.49$\pm$14.80 60.13$\pm$10.23 59.23$\pm$10.40 51.94$\pm$6.06 64.32$\pm$15.92 84.95$\pm$8.94 Cov+EN Ins 71.01$\pm$14.33 59.34$\pm$10.89 59.37$\pm$9.86 52.06$\pm$5.39 64.67$\pm$14.79 84.53$\pm$8.81 FgMDM Ins 71.61$\pm$13.86 60.64$\pm$10.24 59.45$\pm$10.29 52.53$\pm$5.32 65.09$\pm$14.73 85.17$\pm$8.36 Phase-SPDNet Ins 71.84$\pm$16.75 68.83$\pm$17.41 62.04$\pm$12.02 65.37$\pm$12.91 68.31$\pm$14.17 87.03$\pm$6.13
In this study, the situation is even rosier: Phase-SPDNet , used with either imaginary or instantaneous coherence features, emerges as the best pipeline across all datasets. When utilizing imaginary coherence, SPDNet alone already achieves statistically superior results compared to state-of-the-art approaches employing the same features. However, the augmentation procedure introduces a significant performance boost, a finding supported by the meta-effect analysis shown in Fig A1 (c).
In the case of instantaneous coherence, on the other hand, the impact of the augmentation procedure is even more significant. In fact, with this feature, the standard SPDNet fails to exhibit statistically superior performance in comparison to the state-of-the-art pipelines. Phase-SPDNet , on the other hand, stands out as the top-performing pipeline, showing that the augmentation procedure brings a statistically significant influence on classification outcomes (see Fig A2 (a)).
Appendix B Additional Results
Table A2: Parameters common to all DL pipelines.
Parameter Value
Epoch 300
Batch Size 64
Validation Split 0.1
Loss Sparse Categorical Crossentropy
Optimizer Adam
Learning Rate = 0.001
Callbacks ES Early Stopping
Patience = 75
Monitor = Validation Loss
Callbacks LR ReduceLROnPlateau
Patience = 75
Monitor = Validation Loss
Factor = 0.5
Figure A1: Result for right hand vs left hand classification, using within-session evaluation for state-of-the-art imaginary coherence pipelines. Plot (a) provides the relative improvement of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet against SPDNet, COV+EN, and FgMDM, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.
Figure A2: Result for right hand vs left hand classification, using within-session evaluation for state-of-the-art instantaneous coherence pipelines. Plot (a) provides the relative improvement of the method considered with respect to the standard SPDNet of the different pipelines considered. Plot (b) shows a combined meta-analysis (over all datasets) of the different pipelines. It shows the significance of the algorithm on the y-axis being better than the one on the x-axis. The gray level represents the significance level of the ROC-AUC difference in terms of t-values. We only show significant interactions ($p<0.05$). Plots (c), (d), and (e) show the meta-analysis of Phase-SPDNet against SPDNet, COV+EN, and FgMDM, respectively. We show the standardized mean differences of p-values computed as a one-tailed Wilcoxon signed-rank test for the hypothesis given in the plot title. The gray bar denotes the $95\%$ interval. * stands for $p<0.05$, ** for $p<0.01$, and *** for $p<0.001$.